OpenStudy (toxicsugar22):

hi can u help me solve this 3^(9-3x)=27

OpenStudy (angela210793):

3^(9-3x)=27 3^(9-3x)=3^3 since the base is the same=3 we can equal the exponents as well --> 9-3x=3 -3x=3=9 -3x=-6 3x=6 x=6/3 x=2